# MTU: The Multifunction Tree Unit in zkSpeed for Accelerating HyperPlonk

### v1

@@ 3.1.3 Merkle Tree
3.1.4 Multiplication Tree and Product MLE
@@
Binary tree-style product computations are not limited to SumCheck protocols; many cryptographic algorithms require computing products of the form $\prod_{i=0}^{2^{\mu}-1}f_{i}$. In cryptographic contexts, these products typically involve modular multiplication, which is significantly more expensive than standard integer multiplication. A simple way is to accumulate the product one by one, which will incur a high latency because of the data dependency. Therefore, it is crucial to optimize such computations using binary tree structures to reduce computational cost.
@@
One notable example is the Product MLE computation in HyperPlonk, which resembles a standard multiplication tree. However, unlike a typical tree that only outputs the final product at the root, the Product MLE computation requires the outputs from all intermediate levels of the tree. These intermediate values are essential for downstream steps in the protocol. The multiplication tree and Product MLE workloads follow a similar structural pattern to that shown in Figure 2, with the primary difference in the node operation. Instead of modular addition and multiplication, each simply computes a modular multiplication of two nodes from the previous level. The Product MLE workload requires outputting the intermediate results (levels starting from Level 2), which increases bandwidth demands.
@@
SumCheck. At each round $i$ of the SumCheck protocol, one might observe that the summation $\sum_{x=0}^{2^{i}-1}f(x)$ follows a tree-based reduction pattern. However, unlike modular multiplication, modular addition is relatively inexpensive and incurs minimal latency (e.g., in zkSpeed [13], mod-add has only one stage pipeline). As a result, the accumulation dependency does not significantly impact the throughput or total latency. A simple accumulator is therefore sufficient to compute the sum efficiently, without the need for a full binary reduction tree. The modular multiplication has a ten-stage pipeline, thus the accumulator method would significantly influence the latency.
3.2 General Binary Tree Traversals
@@
Binary tree computations can be traversed in various ways, with the two most common strategies being breadth-first search (BFS, or level-order traversal) and depth-first search (DFS, including inorder, preorder and postorder traversal). In DFS, computation progresses by recursively visiting one branch of the tree down to its leaves before backtracking. In contrast, BFS processes the tree level by level, evaluating all nodes at a given depth before moving to the next.
@@
BFS. Each traversal strategy offers trade-offs. BFS is straightforward to implement and control, making it a natural choice in software for binary tree workloads mentioned above. Its dataflow typically involves reading all nodes at a given level from memory, performing computations in parallel, and then writing results back to memory. In the example of Figure 2, BFS computation follows Level 2, 3 and so on. This regular, level-by-level structure is amenable to parallelism. However, BFS has significant hardware design challenges. If implemented entirely on-chip, it requires storing all intermediate levels of the tree, which can incur prohibitive memory costs. For example, processing Build MLE of $2^{23}$ workload size would require 128 MB of on-chip SRAM, and it grows linearly with the size. This is inefficient for typical ASIC designs. Alternatively, if intermediate levels are stored off-chip, the design becomes bandwidth-limited, and the achievable parallelism is constrained by memory access throughput. These limitations motivate the need for more memory-efficient traversal strategies that balance parallelism with resource constraints. Hardware accelerators like Nocap [31] applied BFS for its Merkle tree computation.
Table 1: Different binary tree traversal methods.
@@
DFS. DFS offers substantial memory and bandwidth savings compared to BFS through on-chip data reuse. Since DFS computes each branch to completion before moving on, intermediate results are consumed immediately and do not need to be stored for long durations. For example in Figure 2, it follows the order $f_{0}$, $f_{1}$ to $f_{16}$, and then $f_{2}$, $f_{3}$, $f_{17}$, $f_{24}$, and so on. Due to the data dependency, DFS is naturally less compatible with parallel hardware execution. A naive approach (i.e., allocating multiple PEs and feeding them in DFS order) will result in complex scheduling and high latency due to deep recursive computation. Initial inputs will be interrupted since the PE prioritizes the process of the deeper level. The arrival rate of the initial inputs becomes irregular, leading to an initiation interval (II) greater than one, where not all PEs can accept new initial input data every cycle.
One way to simplify DFS-based scheduling is to divide the tree into multiple subtrees and assign each subtree to a dedicated PE. In Figure 2, we divide the workload into two subtrees as shown in the red dash-boxes, which enables two parallel PEs to compute each subtree, and their results will be merged at the end. Figure 1 is similar, while the Level 4 will be first prepared and then distributed to parallel PEs to compute each subtree. This can improve concurrency, but since the tree is divided into several subtrees, the initial input or output indices are no longer sequential. If an upstream module (e.g., another accelerator block) is producing tree inputs in strict index order, such DFS traversal cannot be directly pipelined with it, limiting throughput. These challenges highlight the need for traversal strategies that preserve DFS’s reuse efficiency while improving input regularity and pipelining compatibility. Hardware accelerators like UniZK [38] applies similar philosophy of dividing tree, but within each subtree it still performs BFS individually.
@@
These traversal strategies are summarized in Table 1. Importantly, regardless of the traversal method, the asymptotic time complexity remains the same $O(n)$ since each approach must visit all $n$ nodes in the tree once [17]. Ideally, parallelism may improve this latency to $O(n/p)$ for $p$ parallel PEs. The key differences lie in space complexity and their suitability for parallel execution. BFS enables straightforward parallelism across tree levels but incurs a space complexity of $O(n)$ due to the need to store a level of intermediate data. In contrast, DFS has lower memory requirements because intermediate results can be reused and discarded quickly, but it presents greater challenges in coordinating parallel execution and achieving efficient pipelining.
@@
Table 2: Scheduling of MTU DFS accumulator when processing an inverted binary tree. $L4_{0}$ means the index-0 node of Level 4. The PE takes two inputs from previous level, and produce one output as the next level.
@@
Input A - $L4_{0}$ $L4_{2}$ $L4_{4}$ $L5_{0}$ $L4_{6}$ $L4_{8}$ $L5_{2}$ $L4_{10}$ $L6_{0}$ $L4_{12}$ $L5_{4}$ $L4_{14}$ $L4_{16}$ $L5_{6}$ $L4_{18}$ $L6_{2}$ $L4_{20}$ $L5_{8}$ $L4_{22}$ $L7_{0}$ $L4_{24}$ $L5_{10}$ $L4_{26}$
Input B - $L4_{1}$ $L4_{3}$ $L4_{5}$ $L5_{1}$ $L4_{7}$ $L4_{9}$ $L5_{3}$ $L4_{11}$ $L6_{1}$ $L4_{13}$ $L5_{5}$ $L4_{15}$ $L4_{17}$ $L5_{7}$ $L4_{19}$ $L6_{3}$ $L4_{21}$ $L5_{9}$ $L4_{23}$ $L7_{1}$ $L4_{25}$ $L5_{11}$ $L4_{27}$
Output - - $L5_{0}$ $L5_{1}$ $L5_{2}$ $L6_{0}$ $L5_{3}$ $L5_{4}$ $L6_{1}$ $L5_{5}$ $L7_{0}$ $L5_{6}$ $L6_{2}$ $L5_{7}$ $L5_{8}$ $L6_{3}$ $L5_{9}$ $L7_{1}$ $L5_{10}$ $L6_{4}$ $L5_{11}$ $L8_{0}$ $L5_{12}$ $L6_{5}$
Table 3: Scheduling of MTU DFS accumulator when processing a forward binary tree (Build MLE). $L7_{0}$ means the index-0 node of Level 7. The PE takes an input from previous level, and produce two outputs as the next level.
@@
Input $L8_{0}$ $L7_{0}$ $L6_{0}$ $L5_{0}$ $L6_{1}$ $L5_{1}$ $L7_{1}$ $L5_{2}$ $L6_{2}$ $L5_{3}$ $L8_{1}$ $L5_{4}$ $L6_{3}$ $L5_{5}$ $L7_{2}$ $L5_{6}$ $L6_{4}$ $L5_{7}$ $L5_{8}$ $L6_{5}$ $L5_{9}$
Output A - $L7_{0}$ $L6_{0}$ $L5_{0}$ $L4_{0}$ $L5_{2}$ $L4_{2}$ $L6_{2}$ $L4_{4}$ $L5_{4}$ $L4_{6}$ $L7_{2}$ $L4_{8}$ $L5_{6}$ $L4_{10}$ $L6_{4}$ $L4_{12}$ $L5_{8}$ $L4_{14}$ $L4_{16}$ $L5_{10}$
Output B - $L7_{1}$ $L6_{1}$ $L5_{1}$ $L4_{1}$ $L5_{3}$ $L4_{3}$ $L6_{3}$ $L4_{5}$ $L5_{5}$ $L4_{7}$ $L7_{3}$ $L4_{9}$ $L5_{7}$ $L4_{11}$ $L6_{5}$ $L4_{13}$ $L5_{9}$ $L4_{15}$ $L4_{17}$ $L5_{11}$
4 Hardware-Friendly Hybrid Traversal
To balance the trade-offs between space complexity and parallelism, MTU introduces a hybrid traversal strategy tailored for hardware efficiency. This method combines the low-memory advantages of DFS with the parallelism-friendly nature of BFS to reduce overall traversal time.
The starting point of our design is rate-matching the upstream. For an inverted binary tree as in Figure 2, if eight nodes of Level 1 arrive per cycle, four PEs will be allocated to process them in parallel, and two PEs are used for the second level, and one PE for the third level. As shown in Figure 3, these seven PEs are connected to form a pipeline, enabling the first to third levels to be grouped and computed in a BFS fashion. That is, nodes in Level 1 to 3 are grouped (as the red dash-boxes in Figure 2), and processed by the seven PEs pipeline. Starting from the fourth level, the generation rate naturally becomes one per cycle, which aligns with the capacity of a single PE, making DFS a suitable choice for deeper levels. Thus, the traversal is divided: the input levels are handled using BFS to match the Level 1 input rate and ensure continuous input indexing. The intermediate data are reused in the pipeline. On the other hand, the deeper levels (starting at Level 4) adopt DFS which has low space complexity. This hybrid scheme maintains parallelism with low space complexity, avoids off-chip bandwidth bottlenecks through on-chip reuse, and simplifies scheduling and control logic. It eliminates the need to store entire intermediate levels, making it practical for large problem sizes. Furthermore, the design is scalable, as the number of PEs can be adjusted to match the input rate, forming a full pipeline with other accelerator units.
Figure 3: An example architecture of the MTU with 8 PEs for Hybrid Traversal. Each PE supports both modular arithmetic and hashing. Their connection enables both forward and inverted binary trees computation. All PEs are interconnected to support pipelined dataflow and can directly output results of the accelerator.
4.1 MTU Architecture
In this section, we present the architecture of MTU. Figure 3 illustrates an example configuration. The MTU is composed of several PEs that are connected to support both standard (forward) and inverted binary tree traversals. In Figure 3, the inverted tree mode (workloads like Merkle tree commitments, multiplication tree) MTU can accept eight inputs per cycle. That is, the PEs will take eight elements of Level 1 (e.g., $f_{0}$ to $f_{7}$ in Figure 2) into the pipeline each cycle, and produce a Level 4 element (e.g., $f_{28}$) into the last PE to work in DFS. The architecture can be extended on demand. Putting more PEs in the front allows the MTU to process more inputs per cycle. Each PE is equipped with a SHA3 engine to accelerate hash-based operations required in Merkle Tree, along with modular addition and multiplication units to support arithmetic workloads (e.g., MLE Evaluation and Multiplication Tree).
Scheduling. We show the scheduling of MTU’s DFS Accumulator PE in Table 2, assuming an MTU configuration as Figure 3 and the PE latency is one cycle. For a large inverted binary tree (e.g., $2^{20}$ nodes in Level 1), since Level 4 elements come at one per cycle, the generation of Level 5 happens every two cycles, as well as the Level 6 happens every four cycles and so on. The temporary result (e.g., $L5_{0}$) will be buffered in a small SRAM and consumed once the input pair is ready to launch. This give us a chance to interleave to schedule levels. The scheduling depends on the “generation rate” of each level. It can be efficiently handle by a simple controller, which prioritizes the computation of a deeper level. If the inputs of a deeper level are ready, it will be triggered to compute. Initially, there are gaps in the scheduling, as it must wait two cycles for each input pair. However, once multiple levels are processed, the gaps will be filled. Each PE is also connected to the output ports (red wires in Figure 3) to support Product MLE computation, which requires all elements of every level.
Figure 4: CPU performance across different workloads using BFS and DFS traversal schemes. Each workload is evaluated at size $2^{20}$.
On the other hand, in the forward binary tree (for operations like Build MLE), the MTU can generate up to eight outputs per cycle. PEs are connected in blue wires in Figure 3. Different from the inverted binary tree, here the four PEs on the left are generating the output level (i.e., Level 1). Nodes in Level 3 to 1 are grouped (as the red dash-boxes in Figure 1), and processed by the seven PEs pipeline. Following the similar DFS scheme, the last PE accumulator is used to generate Level 4 and above. Its scheduling is illustrated in Table 3. Nodes of Level 4 are produced at an average rate of one per cycle, which enables the computation of Level 3 to be launched every cycle. As a result, eight elements of Level 1 can be output at the end in a continuous stream.
@@ 4.1 MTU Architecture
The MTU Hybrid traversal, working with the DFS Accumulator PE, enables recursive computation of deeper levels with minimal memory overhead. Its scheduling follows a fixed pattern, simplifying control logic. The key design insight is to match the generation rates between adjacent levels: nodes in Levels 1 to 3 are processed in groups by the pipeline formed from the seven PEs, while the DFS Accumulator PE works by aligning its input or output rates with the group of seven PEs.
5 Methodology
In this section, we detail our experimental setup for evaluating the MTU architecture.

### v2

@@ 3.1.3. Merkle Tree
Figure 2. Dataflow for computing the MLE evaluation workload (input size $2^{4}$). Final output is marked in blue. Partitioning into two subtrees is shown with red dashed boxes to illustrate parallel DFS processing.
3.1.4. Multiplication Tree and Product MLE
@@
Binary tree-style product computations are not limited to SumCheck protocols; many cryptographic algorithms require computing products of the form $\prod_{i=0}^{2^{\mu}-1}f_{i}$. In cryptographic contexts, these products typically involve modular multiplication, which is significantly more expensive than standard integer multiplication. A simple way is to accumulate the product one by one, which will incur high latency costs because of the data dependency. Therefore, computations using binary tree structures must be optimized to reduce computational cost.
@@
One notable example is the Product MLE computation in HyperPlonk, which resembles a standard multiplication tree. However, unlike a typical tree that only outputs the final product at the root, the Product MLE computation requires the outputs from all intermediate levels of the tree. These intermediate values are essential for downstream steps in the protocol. The multiplication tree and Product MLE workloads follow a similar structural pattern to that shown in Figure 2, with the primary difference in the node operation. Instead of modular addition and multiplication, each simply computes a modular multiplication of two nodes from the previous level. Product MLE’s requirement of outputting the intermediate results (levels starting from Level 2) increases bandwidth demands.
@@
SumCheck. At each round $i$ of the SumCheck protocol, one can observe that the summation $\sum_{x=0}^{2^{i}-1}f(x)$ follows a tree-based reduction pattern. However, unlike modular multiplication, modular addition is relatively inexpensive and incurs minimal latency (e.g., in zkSpeed (Daftardar et al., 2025b), mod-add is a one stage pipeline). As a result, the accumulation dependency does not significantly impact the throughput or total latency. An accumulator is therefore sufficient to compute the sum efficiently, without the need for a binary reduction tree. For product computation, accumulator method would significantly influence the latency because of long mod-mul latency.
3.2. General Binary Tree Traversals
@@
Binary tree computations can be traversed in various ways, with the two most common strategies being breadth-first search (BFS, or level-order traversal) and depth-first search (DFS, including inorder, preorder, postorder traversal). In DFS, computation progresses by recursively visiting one branch of the tree down to its leaves before backtracking. In contrast, BFS processes the tree level by level, evaluating all nodes at a given depth before moving to the next.
Figure 3. MTU architecture with 8 PEs for Hybrid Traversal. Each PE supports modular arithmetic and hashing. Connection enables both forward and inverted binary trees computation. All PEs can directly output results of the accelerator.
@@
BFS (level-ordered). Each traversal strategy offers trade-offs. BFS is straightforward to implement and control, making it a natural choice in software for binary tree workloads mentioned above. Its dataflow typically involves reading all nodes at a given level from memory, performing computations in parallel, and then writing results back to memory. In the example of Figure 2, BFS computation follows Level 2, 3 and so on. This regular, level-by-level structure is amenable to parallelism. However, BFS has significant hardware design challenges. If implemented entirely on-chip, it requires storing all intermediate levels of the tree, which can incur prohibitive memory costs. For example, processing Build MLE of $2^{23}$ workload size would require 128 MB of on-chip SRAM, and it grows linearly with the size. This is inefficient for typical ASIC designs. Alternatively, if intermediate levels are stored off-chip, the design becomes bandwidth-limited, and the achievable parallelism is constrained by memory access throughput. These limitations motivate the need for more memory-efficient traversal strategies that balance parallelism with resource constraints. Hardware accelerators like Nocap (Samardzic et al., 2024) applied BFS for their Merkle tree computations.
DFS. DFS offers substantial memory and bandwidth savings compared to BFS through on-chip data reuse. Since DFS computes each branch to completion before moving on, intermediate results are consumed immediately and do not need to be stored for long durations. For example in Figure 2, it follows the order $f_{0}$, $f_{1}$ to $f_{16}$, and then $f_{2}$, $f_{3}$, $f_{17}$, $f_{24}$, and so on. Due to the data dependency, DFS is naturally less compatible with parallel hardware execution. A naive approach (i.e., feeding instructions in DFS order to multiple PEs) results in complex scheduling and high latency from deep recursion. Since PEs prioritize deeper levels, inputs are stalled, causing irregular arrival rates and an initiation interval (II) greater than one, where not all PEs can accept new inputs every cycle.
Another way to simplify parallelized DFS is to divide the tree into multiple subtrees and assign each subtree to a dedicated PE. In Figure 2, we divide the workload into two subtrees, which enables two parallel PEs to compute each subtree, and their results will be merged at the end. Figure 1 is similar, while Level 4 will be preprocessed. This improves concurrency, but since the tree is divided into several subtrees, the initial input or output indices are no longer sequential. If an upstream module (e.g., another accelerator block) is producing tree inputs in strict index order, such DFS traversal cannot be directly pipelined with it. This can be found in zkSpeed (Daftardar et al., 2025b) where the Fraction MLE feeds the Product MLE unit in the index order. These challenges highlight the need for traversal strategies that preserve DFS’s reuse efficiency while improving input regularity and pipelining compatibility. Hardware accelerators like UniZK (Wang and Gao, 2025) apply a similar philosophy of dividing trees, but within each subtree it still performs BFS individually.
Table 1. Different binary tree traversal methods.
@@
These traversal strategies are summarized in Table 1. Regardless of the traversal method, the asymptotic time complexity remains the same $O(n)$ since each approach must visit all $n$ nodes in the tree once (Erickson, 2023). Ideally, parallelism may improve this latency to $O(n/p)$ for $p$ parallel PEs. The key differences lie in space complexity and their suitability for parallel execution. BFS enables straightforward parallelism across tree levels but incurs a space complexity of $O(n)$ due to the need to store a level of intermediate data. In contrast, DFS has lower memory requirements because intermediate results can be reused and discarded quickly, but it presents greater challenges in coordinating parallel execution and achieving efficient pipelining.
4. Hardware-Friendly Hybrid Traversal
To balance the trade-offs between space complexity and parallelism, MTU introduces a hybrid traversal strategy tailored for hardware efficiency. This method combines the low-memory advantages of DFS with the parallelism-friendly nature of BFS.
The starting point is rate-matching the upstream. For an inverted binary tree in Figure 2, if eight nodes of Level 1 arrive per cycle, four PEs will be allocated to process them in parallel, and two PEs are used for the second level, and one PE for the third level. As shown in Figure 3, these seven PEs are connected to form a pipeline, enabling the first to third levels to be grouped, and processed by the seven PEs pipeline. The pipeline generates one output per cycle at the fourth level, which then can be processed by a single PE in DFS method (Section 4.1.1). Thus, the traversal is divided: Level 1-3 are grouped, handled in level-order to match input rate and ensure continuous indexing. Intermediate data are reused within the pipeline. On the other hand, the deeper levels ($\geq$ Level 4) adopt a depth-first strategy which has low space complexity. This hybrid scheme no longer stores the entire intermediate levels. It balances parallelism and on-chip reuse with simple scheduling. Furthermore, the design is scalable, as the number of PEs can be adjusted to match an input rate, forming a full pipeline with other modules.
4.1. MTU Architecture
Figure 3 illustrates the architecture of MTU. The MTU is composed of several PEs that are connected to support both standard (forward) and inverted binary trees. The inverted tree mode (workloads like Merkle tree commit) MTU can accept eight inputs per cycle. That is, the PEs will take eight elements of Level 1 (e.g., $f_{0}$ to $f_{7}$ in Figure 2) into the pipeline each cycle, and produce a Level 4 element (e.g., $f_{28}$) into the last PE. The architecture can be extended on demand. Putting more PEs in the front allows the MTU to process more inputs per cycle. Each PE is equipped with a SHA3 engine to accelerate hash-based operations required in Merkle Tree, along with modular addition and multiplication units to support arithmetic workloads (e.g., MLE Evaluation and Multiplication Tree).
4.1.1. DFS Accumulator
We show the scheduling of MTU’s DFS Accumulator in Table 2, assuming an MTU configuration as Figure 3 and the PE latency is one cycle. As mentioned above, since Level 4 elements come at one per cycle, Level 5 outputs elements every two cycles, Level 6 outputs elements every four cycles, and so on. The temporary results (e.g., $L5_{0}$) are buffered in a small SRAM and consumed once their corresponding input pairs are ready. This allows interleaving across levels during scheduling. This scheduling is determined by the generation rate of each level. A simple controller can handle this efficiently by prioritizing deeper levels: whenever the inputs of a deeper level are available, that level is triggered to compute. At the beginning, there are unavoidable gaps in the schedule, as each input pair for Level 5 and beyond requires multiple cycles to become ready. However, once multiple levels are processed, these gaps are filled. Each PE is also connected to the output ports (red wires in Figure 3) to support Product MLE computation, which requires all elements of every level.
In the forward binary tree (Build MLE workload), the MTU can generate up to eight outputs per cycle. PEs are connected in blue wires in Figure 3. Unlike the inverted binary tree, here the four leftmost PEs generate the output level (Level 1). Nodes from Level 3 to 1 are grouped and processed by the seven PEs pipeline. Following the similar DFS scheme, the accumulator prepares Level 4 and above, with its schedule shown in Table 3. Level 4 nodes are produced at an average rate of one per cycle, which allows Level 3 computations to launch every cycle and enables a continuous stream of eight Level 1 outputs. The rate of Level 4 determines the scheduling of higher levels. Although there are gaps in the initial cycles, they will be filled as computation proceeds. These gaps are dictated by the launch rates: for example, Level 5 launches every two cycles, while Levels 7 and 8 launch every eight and sixteen cycles, respectively. The scheduler inserts gaps at cycles 2 and 3 to ensure that, after sixteen cycles, the next Level 8 ($L8_{1}$) will not be blocked by overlap.
@@ 4.1.1. DFS Accumulator
Table 2. Scheduling of MTU DFS accumulator when processing an inverted binary tree. $L4_{0}$ means the index-0 node of Level 4. The PE takes two inputs from previous level, and produces one output at the next level.
@@
Input A - $L4_{0}$ - $_{2} - L44L$$_{4}$ $L5_{0}$ $L4_{6}$ - $L4_{8}$ $L5_{2}$ $L4_{10}$ $L6_{0}$ $L4_{12}$ $L5_{4}$ $L4_{14}$ - $L4_{16}$ $L5_{6}$ $L4_{18}$ $L6_{2}$ $L4_{20}$ $L5_{8}$ $L4_{22}$ $L7_{0}$ $L4_{24}$ $L5_{10}$ $L4_{26}$
Input B - $L4_{1}$ - $L4_{3}$ - $L4_{5}$ $L5_{1}$ $L4_{7}$ - $L4_{9}$ $L5_{3}$ $L4_{11}$ $L6_{1}$ $L4_{13}$ $L5_{5}$ $L4_{15}$ - $L4_{17}$ $L5_{7}$ $L4_{19}$ $L6_{3}$ $L4_{21}$ $L5_{9}$ $L4_{23}$ $L7_{1}$ $L4_{25}$ $L5_{11}$ $L4_{27}$
Output - - $L5_{0}$ - $L5_{1}$ - $L5_{2}$ $L6_{0}$ $L5_{3}$ - $L5_{4}$ $L6_{1}$ $L5_{5}$ $L7_{0}$ $L5_{6}$ $L6_{2}$ $L5_{7}$ - $L5_{8}$ $L6_{3}$ $L5_{9}$ $L7_{1}$ $L5_{10}$ $L6_{4}$ $L5_{11}$ $L8_{0}$ $L5_{12}$ $L6_{5}$
Table 3. Scheduling of MTU DFS accumulator when processing a forward binary tree (Build MLE). $L7_{0}$ means the index-0 node of Level 7. The PE takes an input from previous level, and produces two outputs at the next level.
@@
Input $L8_{0}$ - - - $L7_{0}$ - $L6_{0}$ - - $L5_{0}$ $L6_{1}$ $L5_{1}$ $L7_{1}$ $L5_{2}$ $L6_{2}$ $L5_{3}$ $L8_{1}$ $L5_{4}$ $L6_{3}$ $L5_{5}$ $L7_{2}$ $L5_{6}$ $L6_{4}$ $L5_{7}$ - $L5_{8}$ $L6_{5}$ $L5_{9}$
Output A - $L7_{0}$ - - - $L6_{0}$ - $L5_{0}$ - - $L4_{0}$ $L5_{2}$ $L4_{2}$ $L6_{2}$ $L4_{4}$ $L5_{4}$ $L4_{6}$ $L7_{2}$ $L4_{8}$ $L5_{6}$ $L4_{10}$ $L6_{4}$ $L4_{12}$ $L5_{8}$ $L4_{14}$ - $L4_{16}$ $L5_{10}$
Output B - $L7_{1}$ - - - $L6_{1}$ - $L5_{1}$ - - $L4_{1}$ $L5_{3}$ $L4_{3}$ $L6_{3}$ $L4_{5}$ $L5_{5}$ $L4_{7}$ $L7_{3}$ $L4_{9}$ $L5_{7}$ $L4_{11}$ $L6_{5}$ $L4_{13}$ $L5_{9}$ $L4_{15}$ - $L4_{17}$ $L5_{11}$
@@
The MTU Hybrid Traversal, paired with the DFS Accumulator PE, enables deeper recursion with low memory overhead and fixed-pattern control. The design insight is to match the generation rate between levels: grouping Levels 1–3 across seven PEs, while the DFS Accumulator synchronizes with this group to maintain throughput.
5. Methodology